# Convex Approximation of Probabilistic Reachable Sets from Small Samples Using Self-supervised Neural Networks

### v1

@@ Abstract
Probabilistic Reachable Set (PRS) plays a crucial role in many fields of autonomous systems, yet efficiently generating PRS remains a significant challenge. This paper presents a learning approach to generating 2-dimensional PRS for states in a dynamic system. Traditional methods such as Hamilton-Jacobi reachability analysis, Monte Carlo, and Gaussian process classification face significant computational challenges or require detailed dynamics information, limiting their applicability in realistic situations. Existing data-driven methods may lack accuracy. To overcome these limitations, we propose leveraging neural networks, commonly used in imitation learning and computer vision, to imitate expert methods to generate PRS approximations. We trained the neural networks using a multi-label, self-supervised learning approach. We selected the fine-tuned convex approximation method as the expert to create expert PRS. Additionally, we continued sampling from the distribution to obtain a diverse array of sample sets. Given a small sample set, the trained neural networks can replicate the PRS approximation generated by the expert method, while the generation speed is much faster.
Note to PractitionersReachability analysis is one of the popular methods for solving real-time motion planning problems for dynamic systems under uncertainty while ensuring safety. However, determining the reachable set is challenging. Traditional methods are computationally expensive and often require extensive system information. Meanwhile, neural networks have been shown to construct distributions effectively. Hence, we propose leveraging the capability of neural networks to model the distribution of reachable sets. To simplify the creation of training datasets, we incorporate self-supervised learning in training the neural network. The proposed method offers several advantages: it is easy to use, generates accurate results quickly, and requires less system information. However, as neural networks are statistical methods relying on data-driven insights, their effectiveness inherently depends on the availability of sufficient historical data for the system.
Index Terms:
Probabilistic reachable set, Convex approximation, Deep Learning
@@ I Introduction
Autonomous system operations face various uncertainties and hazards [1]. The reachable set approach is an important method in collision avoidance [2, 3], safety verification [4] and path planning [5] for Autonomous systems. However, finding the deterministic reachable sets with 100% confidence level is difficult [6]. Meanwhile, the deterministic reachable sets may be too large and over-conservative to be applicable [7]. Finding the Probabilistic Reachable Set (PRS) is more realistic, incorporating uncertainties and stochastic elements in the system dynamics or initial conditions [8]. A brief visualization of PRS is shown in Fig. 1, where PRS with different confidence levels are colored.
Figure 1: Probabilistic Reachable set represents the set of possible states a dynamical system can reach from an initial state. For instance, a PRS with a 95% confidence level means that 95% of the possible system trajectories, considering uncertainties or noise, will remain within the set bounds
I-A Review
Finding PRS remains challenging. One method for determining reachable sets under bounded uncertainties is the Hamilton-Jacobi (HJ) reachability analysis. This approach formulates the problem as a partial differential equation that can be addressed using numerical techniques. Although this method is manageable in scenarios with fewer dimensions, its efficiency significantly diminishes as the dimensionality increases [9][10]. Additionally, HJ reachability analysis requires detailed information about the dynamics or uncertainties. Meanwhile, strong assumptions are required, such as the absence of time correlations in parameter uncertainties. These limitations make it unsuitable for certain realistic situations [9].
Fortunately, we can identify the reachable sets using a learning-based method given a large dataset from an unknown distribution. Generating a reachable set with machine learning/data-driven has been explored. A trained locally-weighted regression algorithm and nonlinear SVM classifiers can predict the reachability of the two-point boundary value problem (2PBVP) [11]. Other data-driven methods such as the Gauss process [12], Monte Carlo [13], and stochastic disturbance [14] have also been developed to approximate reachable set. However, all of those methods have significant limitations in practice for three aspects: a) they all require a lot of samples to perform; b) they cannot provide fast and accurate probabilistic reachable set with adjustable confidence levels; c) their PRS bound is not convex and linear.
Deep learning technology has seen significant advancements recently in computer vision [15] and nature language process [16]. And those methods are further adopted in many different fields such as finance [17], cyber security [18], healthcare [19, 20] and robotics [21, 22, 23, 24]. Deep learning-based methods have been effectively applied to solve obstacle avoidance [25] and path planning problems [26, 27], demonstrating their suitability for addressing challenges in autonomous control tasks. The Universal Approximation Theorem [28], guarantees that a feedforward neural network can approximate any nonlinear real-valued function. The advantages of neural networks include their ability to generate complex outputs from simple inputs, as seen in generative models. Previous research shows that neural networks can work when massive data is massively missing [29] and can determine uncertainty [30]. Neural networks can rapidly extract the texture information of pavement from just a few images [31]. Neural networks indeed offer significant advantages in terms of reduced computing time. Also, the parallel nature of neural networks can lead to significant reductions in computing time [32]. Furthermore, Transformers [16] and self-supervised learning have revolutionized the field of artificial intelligence [33] and have become essential tools for developing more efficient, and versatile AI models when limited sample is available [34]. Therefore, we propose designing a transformer-based neural network to generate PRS, and we further employ self-supervised learning techniques to train this network.
I-B Contribution
In this paper, we propose a method that uses self-supervised and multi-label training to train a transformer-based neural network for generating convex approximations of PRS from a small number of samples. The numerical experiments prove that the neural network can generate reliable convex approximations when the given samples come from a distribution that has been learned by the network. This effective method relies on just one assumption:
  • Assumption 1: a large number of previous samples have been observed before the predicting task.
This assumption is satisfied in many fields such as autonomous driving [35], flight landing prediction [36], and robotics [37]. Our method offers three unique advantages that set it apart from other methods:
  • Advantage 1: No information about the desired system is needed; only a few samples of the desired distribution are required for PRS generating.
  • Advantage 2: The generation speed is extremely fast, making it suitable for online applications.
  • Advantage 3: All the PRS bound generated are convex and linear.
  • Advantage 4: Very few labeled data are needed for training, which simplifies the training process and makes large-scale training feasible.
II Problem Modeling
The main goal of the paper is to generate a convex approximation that can capture the PRS of an unknown static uncertainty whose data samples are gradually observed.
In this section, we will introduce and discuss the concepts of reachable sets and PRS associated with a dynamic system with uncertain discrete-time characteristics  [38, 39].
Let us consider a general discrete-time dynamic system expressed as
$\boldsymbol{s}_{k+1}=\boldsymbol{f}\left(\boldsymbol{s}_{k},\boldsymbol{u}_{k},\boldsymbol{\theta}_{k},\boldsymbol{w}_{k}\right),$ (1)
where, $\boldsymbol{s}_{k}\in\mathbb{R}^{2}$ denotes the state of the system (we only consider 2d dynamic systems in this paper, we use $s$ instead of $x$ like many other dynamic system research to distinguish x-dimension) at time step $k\in\{1,\ldots,T\}$, $\boldsymbol{u}_{k}\in\mathcal{U}_{k}$ denotes an uncertain control input, $\boldsymbol{\theta}_{k}\in\Theta$ denotes an uncertain parameter, $\boldsymbol{w}_{k}\in\mathbb{W}$ denotes an uncertain disturbance, and $s_{0}\in\mathcal{S}_{0}$ denotes the initial state. Diverging from most previous studies [9], this paper allows the sets $\mathcal{U}_{k}\subset\mathbb{R}^{m}$, $\Theta\subset\mathbb{R}^{p}$, $\mathbb{W}\subset\mathbb{R}^{q}$, and $\mathcal{S}_{0}\subset\mathbb{R}^{2}$ to be unbounded. Taking into account the uncertainties in $\boldsymbol{s}_{k}$, $\boldsymbol{u}_{k}$, $\boldsymbol{\theta}_{k}$, and $\boldsymbol{w}_{k}$, it follows that $\boldsymbol{x}_{k+1}$ becomes a random vector following an unknown probability distribution. Then at time $T$, the reachable set $\mathcal{S}_{T}$ of the dynamic system in Eq. 1 is defined to be
[…]
$\boldsymbol{s}_{0}\in\mathcal{S}_{0},\boldsymbol{u}_{k-1}\in\mathcal{U}_{k-1},\boldsymbol{\theta}_{k-1}\in\Theta,\boldsymbol{w}_{k-1}\in\mathbb{W},$
$k\in\{1,\ldots,T\}\},$
where $\boldsymbol{s}_{0}$ denotes the initial state and $\mathcal{S}_{0}$ denotes the initial set. Observe that because the sets $\mathcal{U}_{k-1}$, $\Theta$, $\mathbb{W}$, and $\mathcal{S}_{0}$ can be unbounded, the resulting reachable set $\mathcal{S}_{T}$ could also be unbounded. This implies that determining a bounded set where the state is definitively contained might not be feasible. Therefore, our goal is to generate a bounded set where the probability of the state being within this set exceeds a certain confidence level, which is called a Probabilistic Reachable Set (PRS). At time $k$, a bounded set $\tilde{\mathcal{S}}_{k}$ can be qualified as a PRS of the dynamic system defined in Eq. 1 at confidence level $\alpha$ if and only if
$\operatorname{Pr}\left(\boldsymbol{s}_{k}\in\tilde{\mathcal{S}}_{k}\right)\geq\alpha.$ (3)
It indicates that the probability of the system reaching the PRS at time $k$ exceeds $\alpha$. For example, if $\alpha=95\%$ and the system runs an infinite number of times, the system will be outside the PRS less than 5% of the time. When $\alpha=100\%$ and a PRS $\tilde{\mathcal{S}}_{k}$ does exist, then the reachable set $\mathcal{S}_{k}$ defined in Eq. 2 belongs to the $\tilde{\mathcal{S}}_{k}$.
The previous expert method [40] can find multiple $\tilde{\mathcal{S}}_{k}$ and all the founded $\tilde{\mathcal{S}}_{k}$ are convex and as tight as possible. We define the set $\tilde{\mathcal{S}}_{k}$ using the positions of the four vertices, represented as $\tilde{\mathcal{V}}_{k}\in\mathbb{R}^{4\times 2}$. The trained neural network should be able to imitate the expert and generate a similar PRS that has a confidence level close to $\alpha$, as shown in Eq. 4,
$\tilde{\mathcal{V}}_{k}=NN(w,\mathbf{s_{k}})$ (4)
where $w$ is the trained weights of the neural network, $\mathbf{s_{k}}=\{s_{k}^{i}\}_{i=1}^{n}$ is a collection of samples from the real distribution of the reachable set with $s_{k}^{i}\sim\boldsymbol{s}_{k}$, and $n$ as the number of the samples. The desired $n$ should be smaller than 100.
In the rest of the paper, we only consider the reachable set for a specific single time step, for example, $k=T$. So to simplify the notation, we use $\mathcal{S}$ to refer the actual reachable set, $\tilde{\mathcal{S}}$ to refer the PRS, $\tilde{\mathcal{V}}$ as the four vertices of the PRS, and $\boldsymbol{s}$ as the state of the system, $\mathbf{s}$ as the collection of samples from the reachable set at desired time point.
III Dataset and Expert Method
(a) Example of 2d-binominal distribution and sample expert convex PRS
(b) Example of fan-shape distribution and sample expert convex PRS
(c) Example of GMR trajectory predictor distribution and sample expert convex PRS
Figure 2: Distribution and convex PRS demo
III-A Target distribution
In this paper, we evaluate the proposed method to determine if it can approximate three different probabilistic distributions, all of which are based on the Gaussian distribution, commonly used in aircraft trajectory modeling [41]. Any sample from the target probabilistic distributions is 2-dimensional data, $\mathbf{s}\in R^{2}$.
III-A1 2d-binomial distribution
2d-binomial distribution [38] forms clusters around two different coordinate points. The sampling plot of the 2d-binominal distribution is shown in Fig.2(a). The sample process includes two steps, first is randomly selecting one of two center points, second is adding Gaussian noise to both the x- and y-coordinates of the chosen center point. Let $\boldsymbol{C}_{1}=(x_{1},y_{1})$ and $\boldsymbol{C}_{2}=(x_{2},y_{2})$ represent the two center points. Let $\mathbf{N}=\begin{pmatrix}N_{x}\\ N_{y}\end{pmatrix}$ be the Gaussian noise added to both coordinates, where $N_{x}\sim\mathcal{N}(0,\sigma^{2})$ and $N_{y}\sim\mathcal{N}(0,\sigma^{2})$ are independent Gaussian random variables with mean 0 and variance $\sigma^{2}$. The sample of the 2d-binomial distribution will be $\boldsymbol{s}=\boldsymbol{C}_{i}+\mathbf{N},i\in\{1,2\}$. The probability density function(PDF) of the 2d-binomial distribution is easy to calculate:
$P(\boldsymbol{s})=\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{1},\sigma^{2}\mathbf{I})+\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{2},\sigma^{2}\mathbf{I}),$ (5)
where, $\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{i},\sigma^{2}\mathbf{I})$ represents the bivariate Gaussian distribution centered at $\boldsymbol{C}_{i}=(x_{i},y_{i})$ with covariance matrix $\sigma^{2}\mathbf{I}$ (identity matrix scaled by $\sigma^{2}$, meaning independent noise with the same variance for $x$ and $y$).
III-A2 Fanshape distribution
The fan-shape distribution[38] represents points sampled from a distribution of motion paths influenced by truncated normal distributions of speed and direction, capturing realistic variations in both parameters. The sampling plot of the fanshape distribution is shown in Fig.2(b).
[…]
The position $\mathbf{s}:=[x_{k},y_{k}]^{\top}$ of the flight at time $k$ is modeled by the following discrete-time dynamic system:
$\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}\\ y_{k-1}\end{bmatrix}+\begin{bmatrix}\cos\theta\\ \sin\theta\end{bmatrix}v\Delta t,$
where $[x_{k-1},y_{k-1}]^{\top}$ is the position of the fight at the previous time step $k-1$. $[x_{k},y_{k}]^{\top}$ is the position of the flight at the time step $k$. $v$ is the speed and $\theta$ is heading angle. Due to the uncertainties in $v$ and $\theta$, the position $\mathbf{s}$ is a random vector that obeys a non-Gaussian distribution, which is referred to as the fanshaped distribution. The PDF of the position is difficult to derive.
III-A3 GMR trajectory predictor
GMR trajectory predictor can regress the possibility distribution of the future position of flight near terminal [42] with multiple Gaussian distributions given the past trajectory. The sampling plot of the GMR trajectory predictor distribution is shown in Fig.2(c). The regressed conditional possibility distribution of the future position of flight is defined by the following equation:
$P(\boldsymbol{s}\mid\boldsymbol{s}^{past})=\sum_{k=1}^{K}\pi_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right),$ (6)
where where $\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right)$ are Gaussian distributions with mean $\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$ and covariance $\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$, $K$ is the number of Gaussians, and, $\boldsymbol{s}^{past}_{k}$ is the input and $\boldsymbol{s}$ is the output, where $\boldsymbol{s}^{past}_{k}$ is the combined past trajectory and future trajectory guide of the ego agent. $\boldsymbol{s}$ is the trajectory prediction. $\pi_{k}\in[0,1]$ are weights that sum up to one. If $\pi_{k}$ is larger, the more likely the $\boldsymbol{s}$ belongs to $\mathcal{N}_{k}$. We use three-time step prediction as the desired distribution.
III-B Expert method

### v2

# Learning Probabilistic Obstacle Spaces from Data-driven Uncertainty using Neural Networks

@@ Abstract
Identifying the obstacle space is crucial for path planning. However, generating an accurate obstacle space remains a significant challenge due to various sources of uncertainty, including motion, behavior, and perception limitations. Even though an autonomous system can operate with an inaccurate obstacle space by being over-conservative and using redundant sensors, a more accurate obstacle space generator can reduce both path planning costs and hardware costs. Existing generation methods that generate high-quality output are all computationally expensive. Traditional methods, such as filtering, sensor fusion and data-driven estimators, face significant computational challenges or require large amounts of data, which limits their applicability in realistic scenarios. In this paper, we propose leveraging neural networks, commonly used in imitation learning, to mimic expert methods for modeling uncertainty and generating confidence regions for obstacle positions, which we refer to as the probabilistic obstacle space. The network is trained using a multi-label, supervised learning approach. We adopt a fine-tuned convex approximation method as the expert to construct training datasets. After training, given only a small number of samples, the neural network can accurately replicate the probabilistic obstacle space while achieving substantially faster generation speed. Moreover, the resulting obstacle space is convex, making it more convenient for subsequent path planning.
@@ I Introduction
Autonomous vehicles operations face various uncertainties and hazards [1]. One of the most important tasks of any autonomous vehicle is to move without collisions, in other words, to remain within a safe space at all times [2]. Safe space is also called safe region [3], obstacle-free space [4] in other works. The first step to find a safe space is usually to find the obstacle space. While many methods exist for generating a safe space under the assumption of a known obstacle space [5], this assumption is often unrealistic in practice, and only a few approaches [6, 7] extend to uncertain obstacles. Meanwhile, many previous path planning algorithms assume the obstacle space is polygonal [8, 9] due to the simplicity of collision checking and the compactness of the representation. The area of the obstacle space should be minimized, as a larger safe space typically results in smoother and shorter trajectories while also reducing the overall planning cost [4]. Building on these insights, this paper proposes a novel method to construct a polygonal probabilistic obstacle space that is both compact and capable of representing uncertain obstacles.
In practice, safety is ensured by declaring a collision whenever the distance between an agent and an obstacle falls below a threshold, even without physical contact [10]. In this work, we focus on obstacles with relatively small volumes, such as flights or UAVs in airspace [11], where the obstacle space is determined primarily by position rather than precise geometry, a common simplification known as the point mass model. An obstacle is certain if a deterministic value represents its position, and uncertain if a probability distribution describes its position. The probabilistic obstacle space is then defined as the confidence region of this distribution.
Uncertainty of an obstacle can be caused by many reasons. The first common reason is perception uncertainty. All of the position-locating sensors, including GPS [12], camera [13], Lidar [14], and radar [15], are subject to measurement noise. It is necessary to construct a probabilistic obstacle space from noisy measurements. Sensor Uncertainty Mitigation (SUM) methods [16] can compute bounds on the errors, but it has a very high computational cost and require diverse and representative data. Meanwhile, those over-conservative approaches make safe spaces very small, making it difficult for planning algorithms to find feasible solutions
Behavioral uncertainty can also render an obstacle uncertain, since different intents lead to entirely different trajectories [17]. Traditional data-driven methods [18, 19] can estimate intent distributions when sufficient historical data are available. However, such distributions are often highly complex and require many parameters to represent, which makes it impractical to directly translate them into obstacle spaces.
Another common source of uncertainty is motion. Even obstacles following a fixed path can exhibit positional variability [6, 20]. For instance, as shown in Fig. 2, a flight may follow its planned trajectory yet still wander around it. Consequently, even with state-of-the-art trajectory predictors, the error between the predicted and actual positions can never be eliminated. The M-estimator of Tukey [21] has been used to estimate obstacle motion from single-camera images, but it is computationally expensive and requires careful parameter tuning.
In this paper, we propose a supervised learning approach that employs multi-label training to train a transformer-based neural network for generating polygonal probabilistic obstacle spaces under uncertainty. Numerical experiments demonstrate that the network can reliably reproduce probabilistic obstacle spaces when the input samples are drawn from obstacle types it has learned. In addition, the proposed method achieves substantially faster generation times compared to baseline approaches. The method offers four key advantages:
  • Advantage 1: The approach does not rely on dynamic models of uncertain obstacles. Instead, it can construct obstacle spaces from only a few noisy measurements or partial observations, reducing the sensing and modeling requirements.
  • Advantage 2: All generated obstacle spaces are convex and linear, which simplifies their integration into path planning algorithms and ensures computational tractability.
  • Advantage 3: Training requires only a small amount of labeled data, which lowers the burden of data collection and annotation and makes the approach practical for large-scale deployment.
  • Advantage 4: The generation process is extremely fast, enabling real-time use in online planning and decision-making systems.
II Problem Statement
The main goal of this paper is to generate the probabilistic obstacle space of an uncertain obstacle from a limited data sample using a neural network:
$o^{\alpha}=NN(w,\mathbf{s}),$ (1)
where $o^{\alpha}$ is the probabilistic obstacle space at confidence level $\alpha\in(0,1)$, $w$ denotes the neural network parameters, and $\mathbf{s}={s^{i}}_{i=1}^{n}$ is the set of observed $n$ samples.
Definition (Probabilistic Obstacle Space). Let $o^{true}$ be the true occupied region of an uncertain obstacle $o$, and let $\mathbb{P}$ denote the underlying probability measure that characterizes uncertainty in the obstacle’s position. A set $o^{\alpha}$ is called a probabilistic obstacle space at confidence level $\alpha$ if it satisfies
$\mathbb{P}[a\in o^{true}\mid a\notin o^{\alpha}]<1-\alpha.$ (2)
That is, if the agent $a$ avoids $o^{\alpha}$, then the probability of entering the true obstacle region $o^{true}$ is bounded above by $1-\alpha$.
Since $o^{true}$ is generally unknown due to imperfect perception and prediction, $o^{\alpha}$ provides a tractable confidence-region approximation. Our definition generalizes several prior notions, including the $\epsilon$-shadow of obstacle spaces [6, 7], Gaussian-distributed obstacles [22], adaptive bounding boxes [23], and $\alpha$-probabilistic occupied sets [24].
The neural network in Eq. (1) is thus trained to approximate the mapping
$f:\mathcal{S}^{n}\to\mathcal{O},\quad f(\mathbf{s})\mapsto o^{\alpha},$ (3)
where $\mathcal{S}^{n}$ is the sample space of $n$ measurements and $\mathcal{O}$ is the family of convex polygonal sets in $\mathbb{R}^{d}$.
We illustrate three representative examples of probabilistic obstacle spaces and their associated data samples.
Figure 1: Motion uncertainty: even with a planned trajectory, the actual obstacle position is affected by disturbances and uncertainties.
Motion uncertainty: Consider the scenario of a moving object following a fixed planned trajectory. However, many factors, such as environmental disturbances, control uncertainties, and human factors, can introduce uncertainty in the actual trajectory. Therefore, as shown in Fig. 1, a single planned trajectory may correspond to infinitely many possible realizations of motion, all of which the ego agent must avoid. To capture this variability, we construct a probabilistic obstacle space that encompasses the possible deviations. Formally, we define the sample set $\mathbf{s}={s^{i}}_{i=1}^{n}$ as the trajectories of $n$ flights following the same planned path.
Figure 2: Behavior uncertainty: flights taking off from the same runway with a different destination will have a different route
Behavior uncertainty: Another important source of uncertainty arises from the intent of other agents. Consider an ego UAV operating in terminal airspace near a busy airport. Although aircraft follow air traffic control clearances and standard procedures, the UAV does not know in advance which clearance each departure will receive. As shown in Fig. 2, multiple aircraft may depart from the same runway within a narrow time window, and each departure may be cleared for a different standard instrument departure depending on destination, weather, and traffic flow. From the UAV’s perspective, the resulting obstacle space is therefore governed by a probability distribution over several candidate routes rather than a single deterministic path. Because such intent distributions are complex and rarely available in closed form, we approximate them by constructing the probabilistic obstacle space from $\mathbf{s}$, defined as the trajectories of $n$ flights that have previously departed from the same runway.
Sensor uncertainty: Fig. 3 illustrates two adjacent circular obstacles whose positions are estimated using noisy sensor data. Because the measurements are corrupted by noise, the exact obstacle positions cannot be determined without calibration or advanced signal processing techniques. Furthermore, a single measurement outcome may be consistent with multiple possible true obstacle positions. To account for this, we define $\mathbf{s}$ as the set of $n$ noisy sensor measurements, from which a probabilistic obstacle space is constructed.
Figure 3: Sensor uncertainty: the position of two adjacent obstacles whose positions are estimated using noisy sensor measurements is uncertain
Since the obstacle space is inherently probabilistic, its exact boundary may be irregular or complex. For path planning, however, a tractable geometric representation is preferred. In particular, polygonal approximations are widely used because they provide a convex, linear structure that can be efficiently integrated into optimization-based planners. To this end, instead of directly generating the full probabilistic obstacle space, we represent it using a convex polygon defined by its vertices. Specifically, the neural network outputs four vertices that characterize the polygonal approximation of the probabilistic obstacle space:
$\tilde{\mathcal{V}}=NN(w,\mathbf{s}),$ (4)
where $w$ denotes the trained parameters of the neural network and $\tilde{\mathcal{V}}$ is the predicted set of polygon vertices.
III Method
III-A Supervised learning and expert method
Figure 4: Supervised learning: we iteratively pick a few samples from $o^{true}$ as inputs to the neural network and task the network with generating the corresponding probabilistic obstacle space.
Previous research [25] has proven that the transformer can process the visible (unmasked) parts of the data, reconstructing the entire input from the encoded, incomplete data (masked data). Therefore, if we only have a few samples from a distribution, the transformer may be able to reconstruct the entire distribution. To train the neural network, we first require expert solutions generated by an expert method to serve as supervision during the training process.